OpenStudy (drewkatski):

9x + 2 = 6x + 2y Make x the subject of the formula :?

OpenStudy (anonymous):

9x + 2 = 6x + 2y (subtract 6x from both sides) 9x - 6x + 2 = 2y 3x + 2 = 2y (subtract 2 from both sides) 3x = 2y - 2 (divide by 3) x = (2y - 2)/3 or x = 2/3y - 2/3

OpenStudy (anonymous):

any questions ?

OpenStudy (drewkatski):

Nope :D Thanks so much! I have an exam tomorrow and I'm just checking up on anything I'm not sure about!